What is the y-coordinate of the vertex of the parabola that is given by the equation below?
y = -1.5(x - 2)^2 + 7




Found 2 solutions by stanbon, scott8148:
Answer by stanbon(75887)

What is the y-coordinate of the vertex of the parabola that is given by the equation below?
y = -1.5(x - 2)^2 + 7
----
vertex: (2,7)
Note that when x = 2 the equation becomes y = 0 + 7 = 7
---
Ans: y = 7
=============
Cheers,
Stan H.
=============

Answer by scott8148(6628)